# spacing=nonfrench

### v1

@@ 5.3 The Equality $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$
then, for $0\leq s\leq 1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ with comparable norms. The conclusions hold in particular for radial-Lipschitz domains.
Before going further, let us proceed from understanding why the conditions (5.3) are assumed. The Helson-Szegö theorem (see Ch.IV in [16]) says that $\log|\varphi^{\prime}|=b_{1}+Hb_{2}$ with $b_{1},b_{2}\in L^{\infty}(\mathbb{T})$ and $\|b_{2}\|_{\infty}<\pi/2$ if and only if $|\varphi^{\prime}|\in A_{2}$ on $\mathbb{T}$, which is only used to guarantee that $B_{2,2}^{s}(\Gamma)$ is stable by conjugation for all $0\leq s\leq 1$ (see Corollary 5.2). For the proof of the corresponding analytic case: $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$, this condition can thus be ignored. Note that $\text{Arg}\varphi^{\prime}=H\log|\varphi^{\prime}|=Hb_{1}-b_{2}$. The extra assumption that $Hb_{1}\in L^{\infty}(\mathbb{T})$ is used to derive that $\text{Arg}\varphi^{\prime}$ is bounded, geometrically which means that the boundary curve cannot spiral too much. For instance, for the double spiral curve, $\text{Arg}\varphi^{\prime}$ keeps increasing and goes to infinity.
Let $r(\theta):\mathbb{R}\to\mathbb{(}0,+\infty)$ be a $2\pi$-periodic continuous function. Then the curve $\Gamma=\{r(\theta)e^{i\theta}:\;\theta\in[0,2\pi)\}$ is called a starlike Jordan curve with respect to 0 and the domain $\Omega$ bounded by $\Gamma$ is called a starlike domain. Recall that if moreover $r(\theta)$ is a Lipschitz function, i.e., there exists a positive constant $M$ such that $\|r^{\prime}\|_{\infty}\leq M$ then the curve $\Gamma$ is called a radial-Lipschitz curve and $\Omega$ is called a radial-Lipschitz domain. Suppose $\varphi$ is a Riemann mapping from $\mathbb{D}$ onto $\Omega$ fixing 0. It is known that $\Omega$ is starlike with respect to 0 if and only if
@@
where $\alpha=1$. The condition $\varphi^{\prime}(0)\neq 0$ implies that $\log(\varphi(z)/z)$ is continuous on $\overline{\mathbb{D}}$ so that $\log|\varphi(z)/z|\in L^{\infty}$, and $\text{Arg}(\varphi(z)/z)\in L^{\infty}$. The condition (5.4), combined with this, shows that $\text{Arg}\varphi^{\prime}\in L^{\infty}$. A simple computation shows that $\Omega$ is a radial-Lipschitz domain if and only if the condition (5.4) holds where the constant $\alpha<1$ depends on $M$, from which it follows that $\Omega$ is a chord-arc domain (see p.172 in [35]). Notice that
@@
We have that a radial-Lipschitz curve $\Gamma$ satisfies the condition (5.3). Consequently, the radial-Lipschitz domain has all properties of Theorem 5.5.
The following abstract interpolation Theorem of Voigt [43] is the main tool in our proof of Theorem 5.5:
@@
We now start the proof of Theorem 5.5. Clearly, we only need to show $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$ with comparable norms. Recall that Bergman spaces with standard weights $A^{p}_{\alpha}$ are defined as the sets of holomorphic functions $g$ on $\mathbb{D}$ such that $g\in L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$, i.e.,
@@
where $p>0$ and $\alpha>-1$. Here, the assumption that $\alpha>-1$ is essential because the space $L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$ does not contain any holomorphic function other than 0 when $\alpha\leq-1$. The definition of $A^{p}_{\alpha}$ is extended to the case where $\alpha$ is any real number; that is consistent with the traditional definition when $\alpha>-1$ (see Theorem 13 in [47]). From the definition it can be seen that if $\alpha<\beta$ then the strict inclusion $A^{p}_{\alpha}\subset A^{p}_{\beta}$ holds.
For our purpose we in particular mention that the space $A^{2}_{-1}$ is the classical Hardy space $E^{2}(\mathbb{D})$, i.e., that is the set of holomorphic functions $g$ in $\mathbb{D}$ such that
@@
Suppose $\varphi$ is a conformal map from $\mathbb{D}$ onto $\Omega$ fixing 0. For any $f\in\mathcal{HB}_{2,2}^{s}(\Omega)$, $0\leq s<1$, we have seen from (5.2) that
@@
Motivated by equivalence (5.6), we define $\mathcal{HB}_{2,2}^{1}(\Omega)$ as the set of analytic functions $f$ on $\Omega$ satisfying $V_{1}(f)^{\prime}\in A^{2}_{-1}$. It is easy to see that $f$ is an anti-derivative of a function in $E^{2}(\Omega)$. To be more precise, recall that $A^{2}_{-1}=E^{2}(\mathbb{D})$ we have
@@
By this definition, $\mathcal{HB}_{2,2}^{1}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$, and we also have $\mathcal{HB}_{2,2}^{0}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{0}(\Gamma)$ with equivalent norms by Theorem 5.3. Combining this with (5.6), we can see that this boils down to prove
@@
Recall that $(A^{2}_{1},A^{2}_{-1})$ is an interpolation pair and $A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\;0<s<1$. Notice that the boundary curve $\Gamma$ of $\Omega$ is a chord-arc curve which is the image of $\mathbb{T}$ under the bi-Lipschitz mapping $\lambda$, we then conclude by (4.3) that $(\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma))$ is an interpolation pair and $\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma)]_{s}=\text{HB}_{2,2}^{s}(\Gamma)$. Note that $\Omega$ is a chord-arc domain, in particular a Smirnov domain, that is equivalent to that the Hardy space $E^{2}(\Omega)$ (i.e., $\text{HB}_{2,2}^{0}(\Omega)$) coincides with the $L^{2}(\Gamma)$ closure of the set $\check{B}$ of polynomials on $\Gamma$ (see Theorem 10.6 in [15]). Since $\text{HB}_{2,2}^{1}(\Gamma)$ is the set of functions $f\in E^{2}(\Omega)$ such that $f^{\prime}\in E^{2}(\Omega)$, the set $\check{B}$ is also a dense subspace of $\text{HB}_{2,2}^{1}(\Gamma)$. By the inclusion relation $\text{HB}_{2,2}^{1}(\Gamma)\subset\text{HB}_{2,2}^{0}(\Gamma)$, we have that $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$.
@@
converges. From simple computation it follows that $|e^{cz^{2}}|^{2}=e^{2c(s^{2}-t^{2})}$, $|(\varphi^{\prime})^{\frac{3}{2}-z}|^{2}=e^{2t\text{Arg}\varphi^{\prime}}|\varphi^{\prime}|^{3-2s}$ and $|f^{\prime}\circ\varphi|^{2}<C_{1}$ since the diameter of $\Omega$ is finite. Here, the constant $C_{1}$ depends on $f$ and $\Omega$. Then,
@@
Set $A:=\|\text{Arg}\varphi^{\prime}\|_{\infty}$ and pick $c\geq A^{2}/4$ we then have that $e^{-ct^{2}}e^{tA}\leq e$. By the Koebe distortion theorem (2.1), $|\varphi^{\prime}(\zeta)|(1-|\zeta|)$ is comparable to $d(\varphi(\zeta),\;\Gamma)$, that is bounded due to the boundedness of the diameter of $\Omega$. By Hölder inequality,
@@
We can then conclude that the integral (5.8) converges, and moreover it has an upper-bound depending only on $\Omega$ and given $f$. That implies that the linear mapping $T_{z}$ sends $\check{B}$ to $A^{2}_{1}$, and for any $f\in\check{B}$ the function $T_{(\cdot)}f:\;S\to A^{2}_{1}$ is bounded. In particular,
@@
$\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}$
@@
By the continuity of parametrized integral, all functions $S\ni z\mapsto T_{z}f\in A^{2}_{1}$,   $\mathbb{R}\ni t\mapsto T_{it}f\in A^{2}_{1}$, and $\mathbb{R}\ni t\mapsto T_{1+it}f\in A^{2}_{-1}$ are continuous.
@@
holds for some constant $C$ depending only on $\alpha$ and any $\zeta\in\mathbb{D}$ (see Theorem 2.1 in [46], p.73 in [47]). For each $\zeta\in\mathbb{D}$, define $l_{\zeta}(g)=g(\zeta)$ for $g\in A^{2}_{1}$. By taking $p=2,\alpha=1$ in (5.9), we see that $\|l_{\zeta}\|\leq C(1-|\zeta|^{2})^{-3/2}$, which implies that $l_{\zeta}\in(A^{2}_{1})^{*}$. Set $A:=\{l_{\zeta};\;\zeta\in\mathbb{D}\}$. Clearly, $A$ is a total subset of $(A^{2}_{1})^{*}$. Now for each $l_{\zeta}\in A$, one can see that the function
@@
Since $\check{B}$ is dense in $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ and $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$ for all $s\in[0,1]$ (see [43]), it follows immediately that $\check{B}$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$. Thus, (5.10) holds true for all $f\in\text{HB}_{2,2}^{s}(\Gamma)$. The bounded inverse theorem implies the inverse of the operator $T_{s}$ is also bounded. This completes the proof of Theorem 5.5.
5.4 The Equality $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$
The aim of this part is to prove the following

### v2

@@ 5.3 The equality $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$
then, for $0\leq s\leq 1$, $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$ with comparable norms. The conclusions hold in particular for the radial-Lipschitz domains.
Before going further, let us start by understanding why conditions (5.3) are assumed. The Helson-Szegö theorem (see Ch.IV in [16]) says that $\log|\varphi^{\prime}|=b_{1}+Hb_{2}$ with $b_{1},b_{2}\in L^{\infty}(\mathbb{T})$ and $\|b_{2}\|_{\infty}<\pi/2$ if and only if $|\varphi^{\prime}|\in A_{2}$ on $\mathbb{T}$, which is only used to guarantee that $B_{2,2}^{s}(\Gamma)$ is stable by conjugation for all $0\leq s\leq 1$ (see Corollary 5.2). Note that $\text{Arg}\varphi^{\prime}=H\log|\varphi^{\prime}|=Hb_{1}-b_{2}$. The extra assumption that $Hb_{1}\in L^{\infty}(\mathbb{T})$ is used to derive that $\text{Arg}\varphi^{\prime}$ is bounded, geometrically which means that the boundary curve cannot spiral too much. For example, for the double spiral curve, $\text{Arg}\varphi^{\prime}$ keeps increasing and goes to infinity.
Let $r(\theta):\mathbb{R}\to\mathbb{(}0,+\infty)$ be a $2\pi$-periodic continuous function. Then the curve $\Gamma=\{r(\theta)e^{i\theta}:\;\theta\in[0,2\pi)\}$ is called a starlike Jordan curve with respect to 0 and the domain $\Omega$ bounded by $\Gamma$ is called a starlike domain. Recall that if, moreover, $r(\theta)$ is a Lipschitz function, i.e., there exists a positive constant $C(\Gamma)$ such that $\|r^{\prime}\|_{\infty}\leq C(\Gamma)$ then the curve $\Gamma$ is called a radial-Lipschitz curve and $\Omega$ is called a radial-Lipschitz domain. Suppose $\varphi$ is a Riemann mapping from $\mathbb{D}$ to $\Omega$ that fixes 0. It is known that $\Omega$ is starlike with respect to 0 if and only if
@@
where $\alpha=1$. The condition $\varphi^{\prime}(0)\neq 0$ implies that $\log(\varphi(z)/z)$ is continuous on $\overline{\mathbb{D}}$ so that $\log|\varphi(z)/z|\in L^{\infty}$, and $\text{Arg}(\varphi(z)/z)\in L^{\infty}$. The condition (5.4), combined with this, shows that $\text{Arg}\varphi^{\prime}\in L^{\infty}$. A simple computation shows that $\Omega$ is a radial-Lipschitz domain if and only if condition (5.4) holds where the constant $\alpha<1$ depends on $C(\Gamma)$, from which it follows that $\Omega$ is a chord-arc domain (see p.172 in [39]). Notice that
@@
We have that a radial-Lipschitz curve $\Gamma$ satisfies the condition (5.3). Consequently, the radial-Lipschitz domain has all the properties of Theorem 5.5.
The following abstract interpolation Theorem of Voigt [49] is the main tool in our proof of Theorem 5.5 and subsequent Theorem 5.7:
@@
We now start the proof of Theorem 5.5. Clearly, we only need to show $\mathcal{HB}_{2,2}^{s}(\Omega)=\text{HB}_{2,2}^{s}(\Omega)$ with comparable norms. Recall that Bergman spaces with standard weights $A^{p}_{\alpha}$ are defined as the sets of holomorphic functions $g$ in the unit disk $\mathbb{D}$ such that $g\in L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$, i.e.,
@@
where $p>0$ and $\alpha>-1$. Here, the assumption that $\alpha>-1$ is essential because the space $L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$ does not contain any holomorphic function other than 0 when $\alpha\leq-1$. The definition of $A^{p}_{\alpha}$ is extended to the case where $\alpha$ is any real number; that is consistent with the traditional definition of $\alpha>-1$ (see Theorem 13 in [53]). From the definition it can be seen that if $\alpha<\beta$ then the strict inclusion $A^{p}_{\alpha}\subset A^{p}_{\beta}$ is satisfied.
For our purpose, we in particular mention that the space $A^{2}_{-1}$ is the classical Hardy space $E^{2}(\mathbb{D})$, that is, the set of holomorphic functions $g$ in $\mathbb{D}$ such that
@@
Suppose that $\varphi$ is a conformal map from $\mathbb{D}$ onto $\Omega$ fixing 0. For any $f\in\mathcal{HB}_{2,2}^{s}(\Omega)$, $0\leq s<1$, we have seen from (5.2) that
@@
Motivated by equivalence (5.6), we define $\mathcal{HB}_{2,2}^{1}(\Omega)$ as the set of analytic functions $f$ in $\Omega$ satisfying $V_{1}(f)^{\prime}\in A^{2}_{-1}$. It is easy to see that $f$ is an anti-derivative of a function in $E^{2}(\Omega)$. To be more precise, recall that $A^{2}_{-1}=E^{2}(\mathbb{D})$ we have
@@
By this definition, $\mathcal{HB}_{2,2}^{1}(\Omega)=\text{HB}_{2,2}^{1}(\Omega)$. Recall that we also have $\mathcal{HB}_{2,2}^{0}(\Omega)=\text{HB}_{2,2}^{0}(\Omega)$ with equivalent norms by Theorem 5.3. Combining this with (5.6), we can see that this boils down to proving
@@
Recall that $(A^{2}_{1},A^{2}_{-1})$ is an interpolation pair, and $A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\;0<s<1$. Notice that the boundary curve $\Gamma$ of $\Omega$ is a chord-arc curve which is the image of $\mathbb{T}$ under the bi-Lipschitz mapping $\lambda$, we then conclude by (4.3) that $(\text{HB}_{2,2}^{0}(\Omega),\text{HB}_{2,2}^{1}(\Omega))$ is an interpolation pair and $\text{HB}_{2,2}^{0}(\Omega),\text{HB}_{2,2}^{1}(\Omega)]_{s}=\text{HB}_{2,2}^{s}(\Omega)$. Note that $\Omega$ is a chord-arc domain, in particular a Smirnov domain, that is equivalent to that the Hardy space $E^{2}(\Omega)$ (i.e., $\text{HB}_{2,2}^{0}(\Omega)$) coincides with the $L^{2}(\Gamma)$ closure of the set $\check{B}$ of polynomials on $\Gamma$ (see Theorem 10.6 in [15]). Since $\text{HB}_{2,2}^{1}(\Omega)$ is the set of functions $f\in E^{2}(\Omega)$ such that $f^{\prime}\in E^{2}(\Omega)$, the set $\check{B}$ is also a dense subspace of $\text{HB}_{2,2}^{1}(\Omega)$. By the inclusion relation $\text{HB}_{2,2}^{1}(\Omega)\subset\text{HB}_{2,2}^{0}(\Omega)$, we have $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)=\text{HB}_{2,2}^{1}(\Omega)$.
@@
converges. From simple computation it follows that $|e^{cz^{2}}|^{2}=e^{2c(s^{2}-t^{2})}$, $|(\varphi^{\prime})^{\frac{3}{2}-z}|^{2}=e^{2t\text{Arg}\varphi^{\prime}}|\varphi^{\prime}|^{3-2s}$ and $|f^{\prime}\circ\varphi|^{2}<C_{1}$ since the diameter of $\Omega$ is finite. Here, the constant $C_{1}$ depends on $f$ and $\Omega$. Then
@@
Set $A:=\|\text{Arg}\varphi^{\prime}\|_{\infty}$ and pick $c\geq A^{2}/4$ we then have that $e^{-ct^{2}}e^{tA}\leq e$. By the Koebe distortion theorem (see (2.1)), $|\varphi^{\prime}(\zeta)|(1-|\zeta|)$ is comparable to $d(\varphi(\zeta),\;\Gamma)$, which is bounded due to the boundedness of the diameter of $\Omega$. By Hölder inequality,
@@
We can then conclude that the integral (5.8) converges, and, moreover, it has an upper-bound depending only on $\Omega$ and given $f$. That implies that the linear mapping $T_{z}$ sends $\check{B}$ to $A^{2}_{1}$, and for any $f\in\check{B}$ the function $T_{(\cdot)}f:\;S\to A^{2}_{1}$ is bounded. In particular,
@@
$\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}.$
@@
By the continuity of the parametrized integral, all functions $S\ni z\mapsto T_{z}f\in A^{2}_{1}$,   $\mathbb{R}\ni t\mapsto T_{it}f\in A^{2}_{1}$, and $\mathbb{R}\ni t\mapsto T_{1+it}f\in A^{2}_{-1}$ are continuous.
@@
holds for some constant $C$ depending only on $\alpha$ and any $\zeta\in\mathbb{D}$ (see Theorem 2.1 in [52], p.73 in [53]). For each $\zeta\in\mathbb{D}$, define $l_{\zeta}(g)=g(\zeta)$ for $g\in A^{2}_{1}$. Taking $p=2,\alpha=1$ in (5.9), we see that $\|l_{\zeta}\|\leq C(1-|\zeta|^{2})^{-3/2}$, which implies that $l_{\zeta}\in(A^{2}_{1})^{*}$. Set $A:=\{l_{\zeta};\;\zeta\in\mathbb{D}\}$. Clearly, $A$ is a total subset of $(A^{2}_{1})^{*}$. Now for each $l_{\zeta}\in A$, one can see that the function
@@
Since $\check{B}$ is dense in $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)$ and $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)$ is dense in $\text{HB}_{2,2}^{s}(\Omega)$ for all $s\in[0,1]$ (see [49]), it immediately follows that $\check{B}$ is dense in $\text{HB}_{2,2}^{s}(\Omega)$. Thus, (5.10) holds for all $f\in\text{HB}_{2,2}^{s}(\Omega)$. The bounded inverse theorem implies that the inverse of the operator $T_{s}$ is also bounded. This completes the proof of Theorem 5.5.
5.4 The equality $\mathcal{B}_{p,p}^{s}(\Omega)=B_{p,p}^{s}(\Gamma)$
In this subsection, our goal is to prove the following